Determine whether each of these functions is a bijection from R to R.
f (x) = (x2 + 1)/(x2 + 2)
1
Expert's answer
2021-09-09T13:06:01-0400

"f(x) = \frac{x^2+1}{x^2+2}"

For x=-1 => "f(x) = \frac{1+1}{1+2}= \frac{2}{3}"

For x=1 => "f(x) = \frac{1+1}{1+2}= \frac{2}{3}"

=> Not one-one function

=> Not a bijective (no inverse exists)